OpenStudy (anonymous):

Does the following series converge or diverge? ∑n=1∞[(n√)÷(n2+1)]

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}[\left( \sqrt{n} \right)\div \left( n ^{2} + 1 \right)\]

OpenStudy (mathteacher1729):

Step 1: stop everything and download this flowchart: http://www.math.psu.edu/files/141seriesflowchart.pdf It is from this page(full of more awesome resources) http://www.math.psu.edu/ug/courses/math141

OpenStudy (anonymous):

not really helping

OpenStudy (mathteacher1729):

Following the flowchart, we'd land at the comparison test. Try comparing \[\sum\frac{\sqrt{n}}{n^2+1} \text{less than or equal to ? }\sum\frac{\sqrt{n}}{n^2}\]

OpenStudy (anonymous):

its less than

OpenStudy (anonymous):

how do we know what it converges to?